vgalleg5:

A diameter of a circle has endpoints P(–10, –2) and Q(4, 6). a.Find the center of the circle. b.Find the radius. If your answer is not an integer, express it in radical form. c.Write an equation for the circle.

Vocaloid:

a) the center of a circle is the middle point of its diameter therefore you can use the midpoint formula on the endpoints  let me know if you're having trouble or want to check your solution ^^

vgalleg5:

(11,5)=16?

Vocaloid:

hm, check your calculations again, since we want a point your final answer will be a set of (x,y) values (-10+4)/2 = ? (-2+6)/2 = ?

vgalleg5:

-8,1

Vocaloid:

be careful with parentheses -10 + 4 = -6, divide that by 2 to get -3 -2 + 6 = 4, divide that by 2 to get 2 so (-3,2) is the midpoint and thus the center solution for a)

Vocaloid:

for b) you will take the midpoint (-3,2) from part a) and one of the endpoints (doesn't matter which one, let's use (4,6) and use the distance formula  this can be kind of tricky so let me know if you're having trouble ~

vgalleg5:

Thank you this is awesome, I wish I found this site a long time ago.

vgalleg5:

8.06

Vocaloid:

good but they want it in radical form, so sqrt(65) would be your solution for b) for c) (x-h)^2 + (y-k)^2 = r^2 let (h,k) = (-3,2) [your center] and r = the radius from part b simply plug these values into the equation and leave everything else alone, then you'll have your solution

vgalleg5:

Got it!